Doggie Delights is planning to purchase a new commercial oven for $125,000. They plan to finance the oven through a loan at 6.5% APR compounded bi-weekly (every 2 weeks) over 12 years a. What is the bi-weekly loan payment? b. How much interest will be paid over the 12 years of the loan? c. How much of the first payment is used to pay down the loan, rather than pay interest? d. Doggie Delights is considering upgrading the machine after 7 years. How much will still be owing after 7 years? e. Doggie Delights has decided to pay $650 bi-weekly to pay the loan off faster. How many years will it take to pay off the loan? Thank u!
Answer #1

We use excel formuale to solve this
a)pmt(rate,nper,pv,fv,type)
rate=6.5%/26(since biweekly payments)
nper=12*26 periods
pv=-125000
fv=0
type=0
=pmt(6.5%/26,12*26,-125000,0,0)
=577.48
b)Total interest paid=total amount paid-loan principal
=(26*577.48)-125000
=55172.67
c)interest on first payment=125000*(6.5%/26)=312.50
loan=577.48-312.5=264.98
d)the principal left can be found using cumprinc formuale in excel
=cumprinc(rate,nper,pv,start,end,type)
=-CUMPRINC(6.5%/26,26*12,125000,1,7*26,0)
=60974.13
balance left=125000-60974.13=64025.87
e) use nper formuale in excel
=nper(rate,pmt,pv,fv,type)
=nper(6.5%/26,-650,125000,0,0)
=263 periods
years it is 263/26=10.1 years